1. A club consist of 20 members. In how many ways can a committee of 4 be selected?
Answer by Alan3354(69443)

A club consist of 20 members. In how many ways can a committee of 4 be selected?
-----------------
The 1st choice is 1 of 20, then 1/19, 1/18, 1/17.
That's 20*19*18*17 = 116280 possibilities.
---------------
But, since choosing A,B,C&D is the same as choosing D,C,A & B, it has to be divided by 4*3*2*1 = 24
116280/24 = 4845